# Sparse Attention Post-Training for Mechanistic Interpretability

### v1

@@ Abstract
We introduce a simple post-training method that makes transformer attention sparse without sacrificing performance. Applying a flexible sparsity regularisation under a constrained-loss objective, we show on models up to 1B parameters that it is possible to retain the original pretraining loss while reducing attention connectivity to $\approx 0.3\%$ of its edges. Unlike sparse-attention methods designed for computational efficiency, our approach leverages sparsity as a structural prior: it preserves capability while exposing a more organized and interpretable connectivity pattern. We find that this local sparsity cascades into global circuit simplification: task-specific circuits involve far fewer components (attention heads and MLPs) with up to 100× fewer edges connecting them. These results demonstrate that transformer attention can be made orders of magnitude sparser, suggesting that much of its computation is redundant and that sparsity may serve as a guiding principle for more structured and interpretable models.
@@ 1 Introduction
Scaling has driven major advances in artificial intelligence, with ever-larger models trained on internet-scale datasets achieving remarkable capabilities across domains. Large language models (LLMs) now underpin applications from text generation to question answering, yet their increasing complexity renders their internal mechanisms largely opaque [Bommasani2021FoundationModels]. Methods of mechanistic interpretability have been developed to address this gap by reverse-engineering neural networks to uncover how internal components implement specific computations and behaviors. Recent advances in this area have successfully identified interpretable circuits, features, and algorithms within LLMs [nanda2023progress, olsson2022context], showing that large complex models can, in part, be understood mechanistically, opening avenues for improving transparency, reliability, and alignment [bereska2024mechanistic].
@@
However, interpretability is bottlenecked by the model itself: even with sophisticated reverse-engineering techniques that can faithfully reveal internal algorithms, the underlying computations implemented by large models can still remain highly complex and uninterpretable. Circuits for seemingly simple tasks may span hundreds of interacting attention heads and MLPs with densely intertwined contributions across layers [conmy2023towards], and features can influence each other along combinatorially many attention-mediated paths, complicating attention attribution [kamath2025tracing]. To exemplify this, Figure 1 (top) illustrates the attention patterns of a small, single-head transformer trained solely on a simple two-digit addition task. Here, the model has learned to solve the task in a highly diffused manner, where information about each token is dispersed across all token locations, rendering the interpretation of the underlying algorithm extremely difficult even in this simple case.
@@
The crux of the problem is that models are not incentivised to employ simple algorithms during training. In the present work, we advocate for directly embedding interpretability constraints into model design in a way that induces simple circuits while preserving performance. We focus our analysis on attention mechanisms and investigate sparsity regularisation on attention patterns, originally proposed in [lei2025spartan], as an inductive bias. To demonstrate how sparse attention patterns can give rise to interpretable circuits, we return to the two-digit addition example: Figure 1 (bottom) shows the attention patterns induced by penalizing attention edges during training. Here, the sparsity inductive bias forces the model to solve the problem with much smaller, intrinsically interpretable computation circuits.
Note that we apply this sparsity regularisation scheme as a post-training strategy for pre-trained LLMs. Because the method operates entirely post-training, it can be applied to existing models without re-running pretraining, offering a flexible way to sparsify their attention patterns and improving their interpretabiliy. We show, on models of up to 1B parameters, that the proposed procedure preserves the original cross-entropy loss on pretraining data while reducing the effective attention map to $\approx 0.3\%$ of its edges. We further demonstrate that this intervention yields more interpretable circuit structure. Using standard activation patching and circuit-discovery methods, we find that the sparsified models exhibit dramatically simplified computational graphs, reducing by up to two orders of magnitude the number of edges between components, while retaining essential algorithmic behaviors. Taken together, these results position attention sparsity as a practical inductive tool for surfacing the minimal functional backbone underlying model behavior.
@@ 2.1 Sparse Attention
As self-attention is a key component of the ubiquitous Transformer architecture, a large number of variants of attention mechanisms have been explored in the literature. Related to our approach are sparse attention methods, which are primarily designed to alleviate the quadratic scaling of vanilla self-attention. These methods typically rely on masks based on fixed local and strided patterns [child2019generating] or sliding-window and global attention patterns [beltagy2020longformer, zaheer2020big] to constrain the receptive field of each token. While these approaches are successful in reducing the computational complexity of self-attention, they require hand-defined heuristics that do not reflect the internal computations learned by the model. Besides these fixed-pattern sparse attention methods, Top-$k$ attention, which enforces sparsity by dynamically selecting the $k$ most relevant keys per query based on their attention scores, has also been explored [gupta2021memory]. While Top-$k$ attention enables learnable sparse attention, the need to specify $k$ limits its scope for interpretability for two reasons. First, selecting the optimal $k$ is difficult, and setting $k$ too low can degrade model performance. Second, and more fundamentally, Top-k attention does not allow the model to choose different $k$ for different attention heads based on the context. We argue that this flexibility is crucial for maintaining model performance; indeed, we observe in our experiments (Sec. 2, for example) that our method induces attention heads that attend to a varying number of tokens and lead to emergent circuits that are sparser than even picking $k=1$, as some attention heads can be completely switched off. Instead of using Top-$k$ or hand-defined masks, we employ a sparsity regularisation approach, originally proposed in the context of world modelling [lei2025spartan], which serves as a soft inductive bias that flexibly induces sparse and interpretable circuits.
2.2 Mechanistic Interpretability
@@
Mechanistic interpretability seeks to uncover how internal components of large language models (LLMs) implement specific computations. In Transformers, attention provides a natural handle for analysis, as attention maps reveal directed information flow between tokens and can often be linked to interpretable algorithmic roles [elhage2021mathematical]. Systematic ablation studies [nanda2023progress] provide a first step by measuring performance degradation when specific heads or neurons are removed. Activation patching tests how substituting activations from a source run into a target run affects model outputs [zhang2023towards]. Attribution patching [syed2024attribution] extends this framework by approximating patching effects via local linearisation, enabling scalable estimation of each component’s effective contribution. Together, these approaches allow researchers to isolate sub-circuits, minimal sets of attention heads and MLPs that are causally responsible for a given behavior or task [conmy2023towards]. Attention itself plays a dual role: it both routes information and exposes interpretable relational structure, making it a key substrate for mechanistic study. Our work builds on this foundation by leveraging sparsity to simplify these circuits, amplifying the interpretability of attention-mediated computation while preserving model performance.
While, to the best of our knowledge, our method is the first to demonstrate improved interpretability by regularising the sparsity of self-attention, several complementary works utilise sparsity as a driving principle to render other components of a Transformer interpretable. Focusing on the feed-forward MLPs, works such as Transcoders [dunefsky2024transcoders] and Cross-Layer Transcoders [ameisen2025circuit] have found success in uncovering interpretable structures by approximating MLPs in pretrained models using higher-dimensional MLPs with sparsely activated hidden neurons. Going beyond using sparsity as a post-hoc analysis tool, recent works have investigated sparsity regularisation during training to induce intrinsically interpretable models. [yang2025mixture], for example, replaces MLPs in the model architecture with mixtures of sparsely activated experts composed of interpretable neurons; Sparse Transformer weights [gao2024weightsparse] regularises model weights to enable the discovery of substantially smaller circuits.
@@ 3 Method
    To induce sparse message passing between tokens, we need an attention mechanism that can ‘zero-out’ attention edges, which in turn enables effective L0-regularisation on the attention weights. This is in contrast to the standard softmax attention mechanism, where naive regularisation would result in small but non-zero attention weights that still allow information flow between tokens.
@@
To this end, we leverage the Sparse Transformer architecture in the SPARTAN framework proposed in [lei2025spartan], which uses sparsity-regularised hard attention instead of the standard softmax attention. In the following subsections, we describe the Sparse Transformer architecture and the optimisation setup, highlighting how this approach satisfies the above desiderata.
@@ 3.1 Sparse Attention Layer
$A_{ij}\sim Bern(\sigma(q_{i}^{T}k_{j})),$ (1)
where $Bern(\cdot)$ is the Bernoulli distribution and $\sigma(\cdot)$ is the logistic sigmoid function. This sampling step can be made differentiable via the Gumbel Softmax trick [jang2017categorical]. This binary matrix acts as a mask that controls the information flow across tokens. Next, the message passing step is carried out in the same way as standard softmax attention, with the exception that we mask out the value embeddings using the sampled binary mask,
$SparseAttention(Q,K,V)=\bigg[A\odot softmax(\frac{QK^{T}}{\sqrt{d_{k}}})\bigg]V,$ (2)
@@
Note that during the forward pass, each entry of $A$ is a hard binary sample that zeros out attention edges, which serves as an effective L0 regularisation. Moreover, since the functional form of the sparse attention layer after the hard sampling step is the same as standard softmax attention, pre-trained model weights can be directly used without alterations. [^1] Technically, the sampled $A$ affects the computation. This can be mitigated by adding a bias term inside the sigmoid function to ensure all gates are open at initialisation. Experimentally, we found this to be unnecessary as the models quickly recover their original performance within a small number of gradient steps.
@@ 3.2 Constrained Optimisation
In order to ensure that the models do not lose prediction performance during the post-training procedure, as per desideratum 3, we follow the approach proposed in [lei2025spartan], which employs the GECO algorithm [rezende2018tamingvaes]. Originally developed in the context of regularising VAEs, the GECO algorithm places a constraint on the performance of the model and uses a Lagrangian multiplier to automatically find the right strength of regularisation during training. Concretely, we formulate the learning process as the following optimisation problem,
@@
This can be solved by taking gradient steps on $\theta$ and $\lambda$ alternately. During training, updating $\lambda$ automatically balances the strength of the sparsity regularisation: when $CE$ is lower than the threshold, $\lambda$ decreases, and hence more weight is given to the sparsity regularisation term. This effectively acts as an adaptive schedule which continues to increase the strength of the regularisation until the model performance degrades.
@@ 4 Experiments
(a) Refer to caption
(b) Refer to caption
Figure 2: (a) Cross-entropy loss with respect to the mean proportion of non-zero attention edges across the evaluation dataset. (b) Example of attention patterns in GPT-2 compared with those in a sparse-attention GPT-2 model for a single IOI task instance.
To evaluate the effectiveness of our post-training pipeline, we finetune pre-trained LLMs and compare their prediction performance and interpretability before and after applying sparsity regularisation. In our experiments, two models, GPT-2 [radford2019language] (124M parameters) and LLaMA3-1B [grattafiori2024llama3herdmodels] (1B parameters), are finetuned on the OpenWebText dataset [Gokaslan2019OpenWeb]. Both models are finetuned on sequences of length 64, with the target cross-entropy set at 3.2 (for GPT-2) and 3.5 (for LLaMA), which are the cross-entropy loss achieved on the training set by the respective models. We focus on short context-windows in the current study for computational reasons. In the following evaluations, we aim to answer two key questions: (1) Can sparse models maintain comparable performance to their dense counterparts with sparse message-passing? (2) Do sparse models lead to smaller circuits that are more easily discoverable using standard mechanistic interpretability methods?
4.1 Sparse Attention
We first evaluate the performance and the achieved sparsity of the post-trained models. Figure 2 (left) shows the cross-entropy loss and the mean proportion of non-zero attention edges across the evaluation dataset. Here, both models are able to maintain comparable performance to their dense counterparts, with only a very small number of active attention edges: only 0.2% to 0.3% of the total edges are active on average. To further corroborate that the sparsity post-training effectively rewires the internal computation of LLMs in a non-trivial way, we also evaluate the dense baseline models using top-k attention, where, for each attention head, only the top-k attention weights are kept while the rest are set to zero. The results are shown in Figure 2 (left), which shows that the performance of dense models degrades significantly under top-k attention. This suggests that, in contrast to sparse models, the internal information flow of dense models is diffused across many attention edges. Remarkably, the average number of active edges in the sparse models is significantly lower than in the top-1 attention setting, indicating that some attention heads in the sparse models are not used at all. This is qualitatively demonstrated in Figure 2 (right), which shows an example attention activation for the sparse model, where a large number of attention heads are not attending to anything and hence do not contribute to the computation circuit. In summary, sparse post-training consolidates information flow into a small number of edges while maintaining the same level of performance. Next, we investigate whether these emergent sparse attention patterns yield smaller and more interpretable circuits under standard mechanistic interpretability techniques.
4.2 Interpretability
We focus on two classical circuit-discovery tasks: variants of the Indirect Object Identification (IOI) task and the Greater-Than task. For both tasks, we apply activation patching between correct and corrupted versions of each input. Activation patching yields an importance score for every attention head. These scores can be computed at two levels: (i) a single-sentence level, using a single pair of correct and corrupted inputs, and (ii) a global level, obtained by averaging scores across many task variants. In both settings, the resulting scores induce a ranking over attention heads based on their relative contribution to task performance. We then select the top-$k$ heads under each ranking and intervene on the model by freezing all remaining heads. In the standard dense-attention setting, we freeze non-selected heads to their mean activation, following prior work showing that this preserves model behaviour more faithfully [conmy2023towards]. In contrast, for the sparse-attention model we freeze non-selected heads to zero, which is both natural given the model’s inductive bias and simplifies the intervention procedure. We compare in Figure 6 the results between mean and zero ablation for completion.
(a) Refer to caption
(b) Refer to caption
Figure 3: (a) Logit attribution per sentence keeping only the top-$k$ attention heads selected individually for each sentence via activation patching scores. (b) Logit attribution using a fixed set of top-$k$ attention heads selected globally across all sentences by averaging activation patching scores.
For each value of $k$, we compute the logit difference between the correct and incorrect answers under the intervention, and compare it to the clean model’s logit difference. We analyse the difference of logit differences and normalise the resulting curves by dividing by the absolute value of the first point and inverting the scale, yielding values between 0 and 1. We present normalised curves for both single-sentence top-$k$ head rankings and global top-$k$ rankings. The single-sentence analysis captures the number of heads needed to preserve performance on a specific prompt, while the global analysis reflects the stability and consistency of the underlying circuit across many task instances.

### v2

@@ Abstract
We introduce a simple post-training method that makes transformer attention sparse without sacrificing performance. Applying a flexible sparsity regularisation under a constrained-loss objective, we show on models up to 7B parameters that it is possible to retain the original pretraining loss while reducing attention connectivity to $\approx 0.4\%$ of its edges. Unlike sparse-attention methods designed for computational efficiency, our approach leverages sparsity as a structural prior: it preserves capability while exposing a more organized and interpretable connectivity pattern. We find that this local sparsity cascades into global circuit simplification: task-specific circuits involve far fewer components (attention heads and MLPs) with up to 100× fewer edges connecting them. Additionally, using cross-layer transcoders, we show that sparse attention substantially simplifies attention attribution, enabling a unified view of feature-based and circuit-based perspectives. These results demonstrate that transformer attention can be made orders of magnitude sparser, suggesting that much of its computation is redundant and that sparsity may serve as a guiding principle for more structured and interpretable models.
Machine Learning, ICML
@@ 1 Introduction
Scaling has driven major advances in artificial intelligence, with ever-larger models trained on internet-scale datasets achieving remarkable capabilities across domains. Large language models (LLMs) now underpin applications from text generation to question answering, yet their increasing complexity renders their internal mechanisms largely opaque (Bommasani, 2021). Methods of mechanistic interpretability have been developed to address this gap by reverse-engineering neural networks to uncover how internal components implement specific computations and behaviors. Recent advances in this area have successfully identified interpretable circuits, features, and algorithms within LLMs (Nanda et al., 2023; Olsson et al., 2022), showing that large complex models can, in part, be understood mechanistically, opening avenues for improving transparency, reliability, and alignment (Bereska and Gavves, 2024).
Figure 1: Visualised attention patterns for a 4-layer toy model trained on a simple 2-digit addition task. The main idea of this work is to induce sparse attention between tokens via a post-training procedure that optimizes for attention sparsity while maintaining model performance. In this example, while both models are able to correctly predict the sum, the sparse model solves the problem with a naturally interpretable circuit. Details of this toy setup and more examples are provided in Appendix A
@@
However, interpretability is bottlenecked by the model itself: even with sophisticated reverse-engineering techniques that can faithfully reveal internal algorithms, the underlying computations implemented by large models can still remain highly complex and uninterpretable. Circuits for seemingly simple tasks may span hundreds of interacting attention heads and MLPs with densely intertwined contributions across layers (Conmy et al., 2023), and features can influence each other along combinatorially many attention-mediated paths, complicating attention attribution (Kamath et al., 2025). To exemplify this, Figure 1 (top) illustrates the attention patterns of a small, single-head transformer trained on a simple two-digit addition task. Here, the model has learned to solve the task in a highly diffused manner, where information about each token is dispersed across all token locations, rendering the interpretation of the underlying algorithm extremely difficult even in this simple case.
@@
The crux of the problem is that models are not incentivised to employ simple algorithms during training. In this work, we advocate for directly embedding interpretability constraints into model design in a way that induces simple circuits while preserving performance. We focus our analysis on attention mechanisms and investigate sparsity regularisation on attention patterns, originally proposed in (Lei et al., 2025), as an inductive bias. To demonstrate how sparse attention patterns can give rise to interpretable circuits, we return to the two-digit addition example: Figure 1 (bottom) shows the attention patterns induced by penalising attention edges during training. Here, the sparsity inductive bias forces the model to solve the problem with much smaller, intrinsically interpretable computation circuits.
In this work, we investigate using this sparsity regularisation scheme as a post-training strategy for pre-trained LLMs. We propose a practical method for fine-tuning existing models without re-running pretraining, offering a flexible way to induce sparse attention patterns and enhance interpretability. We show, on models of up to 7B parameters, that our proposed procedure preserves the performance of the base models on pretraining data while reducing the effective attention map to less than $0.5\%$ of its edges. To evaluate our central hypothesis that sparse attention facilitates interpretability, we consider two complementary settings. First, we study circuit discovery, where the objective is to identify the minimal set of components responsible for task performance (Conmy et al., 2023). We find that sparsified models yield substantially simpler computational graphs: the resulting circuits explain model behaviour using up to four times fewer attention heads and up to two orders of magnitude fewer edges. Second, using cross-layer transcoders (Ameisen et al., 2025), we analyse attribution graphs, which capture feature-level interactions across layers. In this setting, sparse attention mitigates the attention attribution problem by making it possible to identify which attention heads give rise to a given edge, owing to the reduced number of components mediating each connection. We argue that this clarity enables a tighter integration of feature-based and circuit-based perspectives, allowing feature interactions to be understood through explicit, tractable circuits. Taken together, these results position attention sparsity as an effective and practical inductive tool for surfacing the minimal functional backbone underlying model behaviour.
@@ 2.1 Sparse Attention
As self-attention is a key component of the ubiquitous Transformer architecture, a large number of variants of attention mechanisms have been explored in the literature. Related to our approach are sparse attention methods, which are primarily designed to alleviate the quadratic scaling of vanilla self-attention. These methods typically rely on masks based on fixed local and strided patterns (Child et al., 2019) or sliding-window and global attention patterns (Beltagy et al., 2020; Zaheer et al., 2020) to constrain the receptive field of each token. While these approaches are successful in reducing the computational complexity of self-attention, they require hand-defined heuristics that do not reflect the internal computations learned by the model.
Beyond these fixed-pattern sparse attention methods, Top-$k$ attention, which enforces sparsity by dynamically selecting the $k$ most relevant keys per query based on their attention scores, has also been explored (Gupta et al., 2021; DeepSeek-AI, 2025). While Top-$k$ attention enables learnable sparse attention, the necessity to specify $k$ limits its scope for interpretability for two reasons. First, selecting the optimal $k$ is difficult, and setting $k$ too low can degrade model performance. Second, and more fundamentally, Top-k attention does not allow the model to choose different $k$ for different attention heads based on the context. We argue that this flexibility is crucial for maintaining model performance.
More recently, gated attention mechanisms (Qiu et al., 2025) provide a scalable and performant framework for inducing sparse attention. In particular, Lei et al. (2025) introduce a sparsity regularisation scheme for world modelling that reveals sparse token dependencies. We adopt this method and examine its role as an inductive bias for interpretability.
2.2 Circuit Discovery
@@
Mechanistic interpretability seeks to uncover how internal components of LLMs implement specific computations. Ablation studies assess performance drops from removing components (Nanda et al., 2023), activation patching measures the effect of substituting activations (Zhang and Nanda, 2023), and attribution patching scales this approach via local linearisation (Syed et al., 2024). Together, these approaches allow researchers to isolate sub-circuits, minimal sets of attention heads and MLPs that are causally responsible for a given behavior or task (Conmy et al., 2023). Attention itself plays a dual role: it both routes information and exposes interpretable relational structure, making it a key substrate for mechanistic study. Our work builds on this foundation by leveraging sparsity to simplify these circuits, amplifying the interpretability of attention-mediated computation while preserving model performance.
2.3 Attribution Graph
Mechanistic interpretability has gradually shifted from an emphasis on explicit circuit discovery towards the analysis of internal representations and features. Recent work on attribution graphs and circuit tracing seeks to reunify these perspectives by approximating MLP outputs as sparse linear combinations of features and computing causal effects along linear paths between them (Dunefsky et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025b). This framework enables the construction of feature-level circuits spanning the computation from input embeddings to final token predictions. Within attribution graphs, edges correspond to direct linear causal relationships between features. However, these relationships are mediated by attention heads that transmit information across token positions. Identifying which attention heads give rise to a particular edge, and understanding why they do so, is essential, as this mechanism forms a fundamental component of the computational graph (Kamath et al., 2025). A key limitation of current attribution-based approaches is that individual causal edges are modulated by dozens of attention components. We show that this leads to feature-to-feature influences that are overly complex, rendering explanations in terms of other features in the graph both computationally expensive and conceptually challenging.
@@ 3 Method
    To induce sparse message passing between tokens, we need an attention mechanism that can ‘zero-out’ attention edges, which in turn enables effective $L_{0}$-regularisation on the attention weights. This is in contrast to the standard softmax attention mechanism, where naive regularisation would result in small but non-zero attention weights that still allow information flow between tokens.
@@
To this end, we leverage the Sparse Transformer architecture in the SPARTAN framework proposed in (Lei et al., 2025), which uses sparsity-regularised hard attention instead of the standard softmax attention. In the following subsections, we describe the Sparse Transformer architecture and the optimisation setup, highlighting how this approach satisfies the above desiderata.
@@ 3.1 Sparse Attention Layer
$A_{ij}\sim\mathrm{Bern}(\sigma(q_{i}^{T}k_{j})),$ (1)
where $\mathrm{Bern}(\cdot)$ is the Bernoulli distribution and $\sigma(\cdot)$ is the logistic sigmoid function. This sampling step can be made differentiable via the Gumbel Softmax trick (Jang et al., 2017). This binary matrix acts as a mask that controls the information flow across tokens. Next, the message passing step is carried out in the same way as standard softmax attention, with the exception that we mask out the value embeddings using the sampled binary mask,
$\mathrm{SparseAttn}(Q,K,V)=\bigg[A\odot\mathrm{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})\bigg]V,$ (2)
@@
Note that during the forward pass, each entry of $A$ is a hard binary sample that zeros out attention edges, which serves as an effective $L_{0}$ regularisation. Moreover, since the functional form of the sparse attention layer after the hard sampling step is the same as standard softmax attention, pre-trained model weights can be directly used without alterations. [^1] Technically, the sampled $A$ affects the computation. This can be mitigated by adding a positive bias term inside the sigmoid function to ensure all gates are open at initialisation. Experimentally, we found this to be unnecessary as the models quickly recover their original performance within a small number of gradient steps.
@@ 3.2 Constrained Optimisation
In order to ensure that the models do not lose prediction performance during the post-training procedure, as per desideratum 3, we follow the approach proposed in (Lei et al., 2025), which employs the GECO algorithm (Rezende and Viola, 2018). Originally developed in the context of regularising VAEs, the GECO algorithm places a constraint on the performance of the model and uses a Lagrangian multiplier to automatically find the right strength of regularisation during training. Concretely, we formulate the learning process as the following optimisation problem,
@@
This can be solved by taking gradient steps on $\theta$ and $\lambda$ alternately. During training, updating $\lambda$ automatically balances the strength of the sparsity regularisation: when $CE$ is lower than the threshold, $\lambda$ decreases, and hence more weight is given to the sparsity regularisation term. This effectively acts as an adaptive schedule which continues to increase the strength of the regularisation until the model performance degrades. Here, the value of $\tau$ is selected as a hyperparameter to ensure that the sparse model’s performance remains within a certain tolerance of the original base model. In practice, the choice of $\tau$ controls a trade off between sparsity and performance: picking a tight $\tau$ can lead to a slower training process, whereas a higher tolerance can substantially speed up training at the cost of potentially harming model performance. In Appendix C, we provide further discussion on this optimisation process and its training dynamics.
3.3 Practical Considerations
One of the main strengths of our proposed method is that, architecturally, the only difference between a sparse Transformer and a normal one lies in how the dot-product attention is computed. As such, most practical training techniques for optimising Transformers can be readily adapted to our setting. In our experiments, we find the following techniques helpful for improving computational efficiency and training stability.
LoRA finetuning (Hu et al., 2022).
Low rank finetuning techniques can significantly reduce the computational requirements for training large models. In our experiments, we verify on a 7B parameter model that LoRA finetuning is sufficiently expressive for inducing sparse attention patterns.
FlashAttention (Dao, 2023)
FlashAttention has become a standard method for reducing the memory footprint of dot-product attention mechanisms. In Appendix B, we discuss how the sampled sparse attention can be implemented in an analogous manner.
Distillation (Gu et al., 2024).
Empirically, we find that adding an auxiliary distillation loss based on the KL divergence between the base model and the sparse model improves training stability and ensures that the behaviour of the model remains unchanged during post-training.
Figure 2: Comparison of model performance between the base OLMo model and the sparsified model evaluated on the various benchmarks. Across all tasks, the performance of the sparse model remains comparable with the base model despite using substantially fewer attention edges.
@@ 4 Experiments
To evaluate the effectiveness of our post-training pipeline, we finetune pre-trained LLMs and compare their prediction performance and interpretability before and after applying sparsity regularisation. We perform full finetuning on a GPT-2 base model (Radford et al., 2019)(124M parameters) on the OpenWebText dataset (Gokaslan and Cohen, 2019). To investigate the generality and scalability of our method, we perform LoRA finetuning on the larger OLMo-7B model (Groeneveld et al., 2024) on the Dolma dataset (Soldaini et al., 2024), which is the dataset on which the base model was trained. The GPT-2 model and the OLMo model are trained on sequences of length 64 and 512, respectively. In the following subsections, we first present a quantitative evaluation of model performance and sparsity after sparse post-training. We then conduct two interpretability studies, using activation patching and attribution graphs, to demonstrate that our method enables the discovery of substantially smaller circuits.
4.1 Model Performance and Sparsity
We begin by evaluating both performance retention and the degree of sparsity achieved by post-training. We set cross-entropy targets of 3.50 for GPT-2 (base model: 3.48) and 2.29 for OLMo (base model: 2.24). After training, the mean cross-entropy loss for both models remains within $\pm 0.01$ of the target, indicating that the dual optimisation scheme effectively enforces a tight performance constraint. To quantify the sparsity achieved by the models, we evaluate them on the validation split of their respective datasets and compute the mean number of non-zero attention edges per attention head. We find that the sparsified GPT-2 model activates, on average, only 0.22% of its attention edges, while the sparsified OLMo model activates 0.44%, indicating substantial sparsification in both cases. Table 1 provides a summary of the results. To further verify that this drastic reduction in message passing between tokens does not substantially alter model behaviour, we evaluate the sparsified OLMo model on a subset of the benchmarks used to assess the original model. As shown in Figure 2, the sparse model largely retains the performance of the base model across a diverse set of tasks. In sum, our results demonstrate that sparse post-training is effective in consolidating information flow into a small number of edges while maintaining a commensurate level of performance.
Model Base CE $\tau$ Final CE Active Attn
GPT-2 3.48 3.50 3.501 0.22%
OLMo 2.24 2.29 2.287 0.44%
Table 1: Performance and sparsity of post-trained models. Final cross-entropy losses closely match the specified targets, while attention sparsity is substantially increased.
4.2 Circuit Discovery with Activation Patching
Figure 3: Attention patterns of the heads required to explain 90% of model behaviour on a copy task. The sparse model requires substantially fewer attention heads. Moreover, the selected heads exhibit the characteristic ‘induction head’ pattern: each token attends to a previous token at a fixed relative offset, effectively copying information forward through the sequence, a pattern well known to implement the copy mechanism in transformer models. Equivalent plots for OLMo can be found in Appendix D.